— Allocating cost to products

Plus Corp. produces fence materials. One division manufactures fence rails and fence posts. As a general rule, more fence rails are produced than posts. For example, during the last week, 900 rails and 30 posts were manufactured. Direct material costs are $6.2 per rail and $6 per post. Direct labor of $400 was attributed to the rail manufacturing operation and $500 was attributed to posts during the past week. Most of the direct labor hour costs are incurred in setting up the automated equipment. In the manufacturing process, it takes about the same amount of time for the equipment to produce one rail as it does to produce one post. This division has $550,000 in annual manufacturing overhead that is allocated based on direct labor costs. The annual direct labor costs are estimated at $88.000. The company produces 250,000 units per year.

1. Prepare a schedule to show the cost assigned to each rail and each post using direct labor costs as the basis for allocating overhead.

2. Prepare a schedule computing the unit costs of rails and posts using units of production as a basis for allocating overhead costs.

3. Explain to management which method of overhead allocation appears more reasonable. Support your answer
Answer #1

Answer

  • Requirement 1

A

Overheads to be allocated

$550,000

B

Amount of direct labor cost as allocation

$88,000

C = A/B

Overheads to be allocated per $ of direct labor cost

$6.25

Overhead allocation:

Working

Rail

Post

A

Amount of Direct labor cost

$400

$500

B

Overheads to be allocated per $ of direct labor cost

$6.25

$6.25

C = A x B

Overheads allocated

$2,500

$3,125

D

No. of units

900

30

E = C/D

Overhead cost per unit

$2.78

$104.17

  • Requirement 2

A

Overheads to be allocated

$550,000

B

No. of units as allocation base

                 250,000

C = A/B

Overheads to be allocated per unit

$2.20

Working

Rail

Post

A

No. of units produced

$900

$30

B

Overheads to be allocated per unit

$2.20

$2.20

C = A x B

Overheads allocated

$1,980

$66

D

No. of units

900

30

E = C/D

Overhead cost per unit

$2.20

$2.20

  • Requirement 3

It is clear from above requirements how different methods will lead to different amount of overheads being allocated.

However the question data states that “it takes about the same amount of time for the equipment to produce one rail as it does to produce one post”.
This means that Allocation based on “Direct Labor cost” should be preferred and is more reasonable.

Answer: Allocation based on Direct Labor Cost [requirement 1] is more reasonable.